# Evaluating Compiler Optimization Impacts on zkVM Performance

### v1

@@ 4.3. RQ3: Comparison to Traditional CPUs
Figure 7 shows the average impact of each optimization pass on zkVM and x86 performance, measured across all benchmarks. For brevity, we exclude passes with an average impact on zkVM execution time, proving time, and x86 execution time that is below 2%. For the majority of passes, the direction of their impact (positive or negative) is similar on both zkVMs and x86. However, the magnitude of the effect is significantly larger on x86. This could be explained by the fact that many LLVM optimizations fail to show their full potential on zkVMs. This is mainly due to a mismatch in the cost model or the underlying heuristics of LLVM, which rely on assumptions that hold only on the traditional architectures. We provide further evidence on this in Section 5.
Figure 8. Number of programs where each optimization pass exhibits: (1) a performance gain on x86 but a loss on RISC Zero, (2) a gain on both x86 and RISC Zero with ¿5% higher gain on x86, (3) a gain on both with ¿5% higher gain on RISC Zero, or (4) a gain on RISC Zero but a loss on x86.
@@ 5. Performance Impact Analysis
Based on our study’s results (Section 4), we examine a selection of passes with interesting performance behaviors. We select these passes based on four criteria: (1) passes that cause considerable performance losses on zkVMs, (2) passes that provide significant performance gains on zkVMs, (3) passes that benefit x86 execution but degrade performance on zkVMs, and (4) passes that improve both x86 and zkVM performance, but with a substantially greater impact on x86.
@@
[…]
Figure 9(a) illustrates a simplified loop extracted from npb-lu where licm incurs this overhead. This comes from the increased number of getelementptr instructions in the LLVM IR, each computing the address of an array element to store the constant 42 (line 5). These extra memory operations in turn cause significant paging pressure with an increase in page-ins and page-outs. Nested loops exacerbate this effect: with a nesting of depth 4 (Figure 9(b)), licm increases cycle count by $2.6\times$, compared to $1.3\times$ for depth 2. Overall, the extra memory pressure stemming from licm is evident when loops iterate over arrays.
Insight 1 (I1): Optimizations that introduce a lot of paging pressure should be avoided. Optimizations that increase paging pressure harm zkVMs. Examples of passes that can potentially increase paging pressure include licm, inline (see our analysis below), and reg2mem. Since page-ins and page-outs are highly expensive in zkVMs, compilers targeting zkVMs must prioritize paging-aware cost models when transforming loops or spilling to memory.
@@
Significant improvements on zkVMs. According to RQ1 (Section 4.1), inline is the most beneficial pass for zkVMs. But is LLVM’s default inlining optimal (Theodoridis et al., 2022)? To find out, we raise the inlining threshold from 225 (default) to 4328 using the -inline-threshold option, allowing the compiler to inline more aggressively even larger functions. Notably, the value of 4328 comes from the OpenTuner’s autotuning results on the NPB and crypto benchmarks (Section 4.2). Compared to -O3 with the default threshold, the higher -inline-threshold improves execution by 6% on RISC Zero and 1% on SP1 on average, with some benchmarks showing dramatic gains, such as +44% on RISC Zero and +40% on SP1 for the npb-bt program. In contrast, increasing the inline threshold degrades x86 performance on average by 1%.
@@
Insight 2 (I2): Increasing the inlining threshold on zkVMs can be very beneficial. A more aggressive inlining policy can yield substantial performance gains on zkVMs because it may reduce call overhead and paging costs. This suggests that compilers targeting zkVMs should: (1) adopt a zkVM-aware cost model for inlining rather than reusing heuristics from traditional CPUs, and (2) tune inlining thresholds per workload, especially for performance-critical code (Section 6.2).
The importance of tuning inlining and using a zkVM-aware model is exemplified by the program in Figure 10 (extracted from the tailcall program). Here, inlining causes major regressions on both zkVM execution ($0.8\times$ speedup) and proving ($0.45\times$ speedup). The root cause is stack spilling: RISC-V has only 32-bit registers, so each 64-bit variable occupies two. After inlining the function work (lines 1 and 12), three u64 variables (outer-loop index i on line 11, inner-loop index j on line 3, and sum on line 2) must coexist. This forces the compiler to spill both j and sum to the stack every inner-loop iteration. This doubles the lw/sw (load/store) instructions, which account for most of the $2\times$ cycle increase. Following I1 (Insight 1), inlining should be avoided when it triggers stack spills, as loads and stores may worsen execution due to expensive page-ins and page-outs in zkVMs (ris, 2025a).
Improvements on x86; degradations on zkVMs. On polybench-durbin, the loop-unroll pass improves x86 execution by 9% but degrades RISC Zero performance by 8.7%. Figure 11 shows an oversimplified version of this benchmark where loop-unroll produces this divergence between x86 and zkVM performance. Overall, x86 sees gains of 9% when loop unrolling this example program compared to the unoptimized baseline, whereas both zkVMs consistently suffer slowdowns of 3%–10% in both execution and proving time.
Table 2. The relative speedup percentage of 4x and 16x unrolling compared to the non-unrolled assembly code.
@@
[…]
4 28.10% 24.25% 6.60% 51.43% 39.51%
16 31.51% 26.42% 1.54% 51.62% 52.66%
When used with -O3 and a high unroll-threshold (aggressive unrolling), loop unrolling becomes beneficial compared with -O3 and no loop unrolling. Specifically, on RISC Zero, execution improves by 16% and proving by 3.7%; on SP1, by 0.5% and 1.7%. However, these gains are still modest compared to x86’s $32\times$ speedup improvement.
To isolate the effect, we manually unroll loops in RISC-V assembly, avoiding interference from other optimizations or vectorization. Table 2 reports speedups for $4\times$ and $16\times$ loop unrolls. This manual effort yields substantial performance gains for both zkVMs, with RISC Zero even surpassing the gains seen on x86. These improvements come mainly from executing fewer loop bookkeeping instructions, such as counter increments, comparisons, and branch jumps.
Insight 3 (I3): Performance gains are observed only when loop unrolling leads to a reduction in cycle count. Since zkVMs cannot exploit instruction-level parallelism, out-of-order, or superscalar execution, performance gains when unrolling loops depend entirely on reducing instruction count. This suggests that compilers targeting zkVMs should adopt a tailored cost model for loop unrolling, prioritizing instruction reduction over traditional hardware assumptions.
@@
Substantial improvement on x86; minimal effect on zkVMs. On polybench-nussinov, simplifycfg improves x86 by 23.6%, but SP1 execution by only 2.6%. Figure 12 shows a minimized function from this benchmark that computes the absolute value of an integer using a branch. simplifycfg replaces the branch with equivalent arithmetic operations (Figures 12(b) and 12(c)). When tested with an array of integers chosen at random, the optimized version runs $2.2\times$ faster on x86, presumably thanks to less branch misprediction overhead. However, on zkVMs, the optimized code increases cycle count by 17.7% on RISC Zero and 7.6% on SP1. Proving time regresses similarly (18.1% and 7.0%), and execution time degrades by 8.5% and 7.7%, respectively.
Insight 4 (I4): branches should not always be eliminated. Compilers targeting zkVMs should be more conservative when removing branches, for two reasons. First, branches are relatively cheap and incur no misprediction penalty compared to traditional architectures. Second, predicated execution increases proving cost, since both paths must be proven.
@@ 6. Implications
We demonstrate the practical value of our study’s results by refining existing LLVM passes to improve zkVM performance and then distill several key lessons from our findings.
@@
Change set 1: Adopting a zkVM-specific cost model. The default RISC-V cost model in LLVM is not zkVM-aware, yet instruction costs on zkVMs differ greatly from traditional CPUs. For example, multiplication costs the same as addition, division is not expensive, while floating-point operations are extremely costly due to emulation (ris, 2025b). Memory access is also far more expensive when it triggers page-ins or page-outs (see Insight 1, Section 5). We address these by updating the RISCVTTIImpl and RISCVTargetLowering classes, which provide target-specific cost models to LLVM’s middle-end so that passes can make better decisions for this architecture. Our modifications encode a zkVM-aware cost model for arithmetic and memory operations (e.g., division is no longer considered expensive).
Change set 2: Updating defaults and heuristics. Many LLVM passes use default parameters and heuristics that are tuned for traditional architectures. This tuning may not be suitable for zkVMs. Adjusting these for zkVM-specific instruction costs, such as making inling agressive (Insight 2) or applying loop unrolling only when it reduces instructions count (Insight 3), can improve generated code performance.
We increase the default inlining threshold and adjust parameters such as inline-call-penalty accordingly. We also update simplifycfg to make branch elimination more conservative and avoiding evaluation of both sides of a predicated operation (Insight 4).
Change set 3: Disabling irrelevant passes. Our RQ1 analysis (Section 4.1) shows that some passes offer little or no benefit on zkVMs, mainly because they rely on hardware features that are unsupported in zkVMs. For example, speculative-execution hoists side-effect-free instructions to reduce branch-misprediction latency on out-of-order, superscalar CPUs. This advantage is irrelevant to zkVMs, which execute RISC-V code strictly in order inside an arithmetic circuit. Similarly, we also disable loop-data-prefetch, and hot-cold-splitting, as they provide no measurable gain on zkVMs.
@@
[…]
The largest regression occurs when proving regex-match on SP1 (+27.3%). This is caused by our changes in the inline’s heuristics, which make SP1 produce 20 proof shards instead of 16. Proof sharding divides the program into smaller proving units; more shards increase proving time due to additional aggregation overhead.

### v2

@@ 4.3. RQ3: Comparison to Traditional CPUs
Figure 7 shows the average impact of each pass on zkVM and x86 performance, measured across all benchmarks, relative to the baseline where no optimizations are applied, including Rust’s MIR optimizations. For brevity, we exclude passes with an average impact on zkVM execution time, proving time, and x86 execution time that is below 2%. For the majority of passes, the direction of their impact (positive or negative) is similar on both zkVMs and x86. For example, inline is the most beneficial pass on both platforms, while licm is the most detrimental pass when applied in isolation. The latter occurs because licm relies on subsequent passes to eliminate redundant getelementptr operations, which cause significant performance regressions (Section 5).
However, the magnitude of the effect is significantly larger on x86. This could be explained by the fact that many LLVM optimizations fail to show their full potential on zkVMs. This is mainly due to a mismatch in the cost model or the underlying heuristics of LLVM, which rely on assumptions that hold only on the traditional architectures. We provide further evidence on this in Section 5.
Figure 8. Number of programs where each pass exhibits: (1) a performance gain on x86 but a loss on RISC Zero, (2) a gain on both x86 and RISC Zero with ¿5% higher gain on x86, (3) a gain on both with ¿5% higher gain on RISC Zero, or (4) a gain on RISC Zero but a loss on x86.
@@ 5. Performance Impact Analysis
We examine the results of our study (Section 4) to distill the key factors that influence zkVM performance and explain why certain optimization passes perform well or poorly.
Figure 9. Impact of representative passes on RISC Zero performance and cost metrics. Each plot shows the percentage change relative to the baseline for proving time and execution time, along with changes in key zkVM cost components: total cycles, dynamic instruction count, and paging cycles. The x-axis represents benchmarks. For execution and proving time, higher values indicate performance improvements, whereas for the cost metrics, higher values correspond to increased overhead.
5.1. Cost Components
According to RQ1 (Section 4.1), inline is the most beneficial pass for zkVMs, while licm is the most detrimental pass for zkVM performance. To understand why, we analyze the underlying cost components. Beyond cycle count, which directly reflects execution cost (Section 4.1), we examine the dynamic instruction count, i.e., the number of executed instructions. Our insight is that since zkVM’s model assigns almost a uniform cost for each instruction, executing fewer instructions could potentially lead to performance gains. Finally, for RISC Zero (the corresponding metric is not available in SP1), we consider paging cycles, which measure the cycles required to move data between the guest program’s memory and the zkVM’s storage system. Page-ins and page-outs introduce high overhead in zkVMs. For example, on RISC Zero a paging operation costs roughly 1,130 cycles (ris, 2025).
Figure 9 illustrates the impact of the two most beneficial passes (inline and always-inline), and the two most detrimental passes (loop-extract and licm) on execution time and proving time in RISC Zero, relative to the baseline with no optimizations. For completeness, we also show the corresponding plots for -O3 and -O0. The other passes and pipelines exhibit similar trends and can be found in our artifact.   [^2] https://github.com/thomasgassmann/zkvm-compiler-optimizations The plots also report the effect of these passes on key zkVM cost components: cycle count, dynamic instruction count, and paging cycles. There is a clear trend: when a pass (e.g., inline) yields significant improvements in execution or proving time, it also leads to substantial reductions in the underlying cost metrics (e.g., inline on polybench-floyd-warshall). Similarly, performance regressions are mainly attributed to tremendous increases in either dynamic instruction count or paging cycles, or all of them (e.g., licm on npb-lu). When performance gains in execution or proving time are more modest, the corresponding reductions in cost metrics are also limited, or different metrics exhibit opposing effects, such as a decrease in executed instructions accompanied by an increase in paging cycles (e.g., always-inline on factorial).
Table 2. Monotonic and linear relationships between zkVM cost metrics and performance, reported as average Kendall’s and Pearson’s coefficients. Correlations are computed per benchmark over optimization variants.
SP1 RISC Zero
Performance metric Cost Metric Kendall’s $\tau$ Pearson Kendall’s $\tau$ Pearson
zkVM exec time Executed instr. 0.87 0.99 0.58 0.97
Paging cycles N/A N/A 0.38 0.88
Total cycles 0.87 0.99 0.59 0.97
Proving time Executed instr. 0.49 0.9 0.5 0.95
Paging cycles N/A N/A 0.33 0.89
Total cycles 0.49 0.89 0.48 0.95
Paging cycles Executed instr. N/A N/A 0.43 0.9
Monotonicity and linearity. To assess whether these cost components can reliably guide compiler optimization decisions for zkVMs, we evaluate the monotonic relationship between dynamic instruction count, paging cycles, and zkVM performance. To this end, for each benchmark, we compute Kendall’s $\tau$ correlation coefficient (KENDALL, 1938) between each cost metric (dynamic instruction count, paging cycles, total cycles) and both zkVM execution time and proving time, across individual passes applied in isolation. We also report the Pearson correlation coefficient, which captures linear relationships. Table 2 summarizes the results.
Overall, dynamic instruction count and paging cycles exhibit moderate-to-strong positive correlations with both execution time and proving time across zkVMs. For example, a Kendall’s $\tau$ value of 0.58 between dynamic instruction count and RISC Zero execution time indicates that, for a given benchmark, there is a 79% probability that an optimization increasing dynamic instruction count also increases execution time. This probability is computed by $(1+\tau)/2$, where $\tau$ is the Kendall’s coefficient. The high Pearson coefficients (>> 0.9) further suggests that this relationship is linear.
For RISC Zero, dynamic instruction count shows stronger correlation with both execution time and proving time than paging cycles. While dynamic instruction count is the primary cost driver, the two metrics are complementary, as reflected by a Kendall’s $\tau$ of 0.43 between them. In 28.5% of cases, the metrics disagree: an optimization may reduce instruction count while increasing paging overhead, or vice versa (e.g., Figure 9; always-inline on factorial).
@@ 5.1. Cost Components
5.2. Case Studies
Based on our study’s results (Section 4), we examine a selection of passes with interesting performance behaviors. We select these passes based on four criteria: (1) passes that cause considerable performance losses on zkVMs, (2) passes that provide significant performance gains on zkVMs, (3) passes that benefit x86 execution but degrade performance on zkVMs, and (4) passes that improve both x86 and zkVM performance, but with a substantially greater impact on x86. Our analysis aims to identify the root causes of these performance behaviors based on the main zkVM cost components (Section 5.1). Then, we derive a set of principles that can guide potential LLVM improvements for better zkVM performance (Section 6.1).
Significant performance regressions on zkVMs. licm is the most detrimental pass for zkVM performance in RQ1 (Section 4.1). The effect is particularly severe on the npb-lu benchmark, where licm slows down proving time by $2.7\times$ on RISC Zero and $1.7\times$ on SP1, and execution by $1.7\times$ on both. The root cause is a sharp increase in paging cycles: +444% on RISC Zero, +69% on SP1. Indeed, across all benchmarks, licm increases paging cycles by 32%.
Figure 10(a) illustrates a simplified loop extracted from npb-lu where licm incurs this overhead. This comes from the increased number of getelementptr instructions in the LLVM IR, each computing the address of an array element to store the constant 42 (line 4). These extra memory operations in turn cause significant paging pressure with an increase in page-ins and page-outs.
@@ 5.2. Case Studies
The pattern is also exemplified by the program in Figure 11 (extracted from the tailcall program). Here, inlining causes major regressions on both zkVM execution ($0.8\times$ speedup) and proving ($0.45\times$ speedup). The root cause is stack spilling: RISC-V has only 32-bit registers, so each 64-bit variable occupies two. After inlining the function work (lines 1 and 10), three u64 variables (outer-loop index i on line 10, inner-loop index j on line 3, and sum on line 2) must coexist. This forces the compiler to spill both j and sum to the stack every inner-loop iteration. This doubles the lw/sw (load/store) instructions, which account for most of the $2\times$ cycle increase. Indeed, the number of the executed instructions increases by 102%, while the number of paging cycles increases by 54%.
Principle 1 (P1): Optimizations that introduce a lot of paging pressure should be avoided, as it harms zkVMs. Examples of passes that can potentially increase paging pressure include licm, inline, and reg2mem. Since page-ins and page-outs are highly expensive in zkVMs, compilers targeting zkVMs must prioritize paging-aware cost models when transforming loops or spilling to memory.
To validate this principle, we manually increase the nesting of the loop of Figure 10(a). Nested loops exacerbate the negative effects of paging pressure: with a nesting of depth 4 (Figure 10(b)), licm increases paging cycles and dynamic instruction count by 46% and 155%, respectively, compared to 7% and 25%, at a depth 2. The extra memory pressure stemming from licm is evident when loops iterate over arrays, as this leads to a higher number of load/store operations.
Significant improvements on zkVMs. According to RQ1 (Section 4.1), inline is the most beneficial pass for zkVMs. But is LLVM’s default inlining optimal (Theodoridis et al., 2022)? To find out, we raise the inlining threshold from 225 (default) to 4328 using the -inline-threshold option, allowing the compiler to inline more aggressively even larger functions. Notably, the value of 4328 comes from the OpenTuner’s autotuning results on the NPB and crypto benchmarks (Section 4.2). Compared to -O3 with the default threshold, the higher -inline-threshold improves execution by 6% on RISC Zero and 1% on SP1 on average, with some benchmarks showing dramatic gains, such as +44% on RISC Zero and +40% on SP1 for the npb-bt program. These performance gains are accompanied by similar decreases in both dynamic instruction count and paging cycles. In contrast, increasing the inline threshold degrades x86 performance on average by 1%.
@@
Principle 2 (P2): Applying inlining selectively based on dynamic instruction count is beneficial for zkVMs. Our analysis in Figure 9 shows that reductions in executed instructions caused by inline often lead to substantial zkVM improvements, primarily by eliminating call overhead. However, as shown in Figure 11, inlining should be avoided if it causes spills to the stack, as load/store instructions increase dynamic instruction count and may trigger expensive memory paging (Figure 11). This suggests that compilers targeting zkVMs should: (1) adopt a zkVM-aware cost model for inlining rather than reusing heuristics from traditional CPUs, and (2) tune inlining thresholds per workload, especially for performance-critical code (Section 6.2).
To validate this principle, we manually inline selected functions in the tailcall benchmark by adding #[inline(always)] annotations, while excluding the function that triggers stack spills. This reduces dynamic instruction count by 5%, yielding a 2% faster execution and eliminating the previous regression.
Improvements on x86; degradations on zkVMs. On polybench-durbin, the loop-unroll pass improves x86 execution by 9% but degrades RISC Zero performance by 8.7%. Figure 12 shows an oversimplified version of this benchmark where loop-unroll produces this divergence between x86 and zkVM performance. Overall, x86 sees gains of 9% when loop unrolling this example program compared to the unoptimized baseline, whereas both zkVMs consistently suffer slowdowns of 3%–10% in both execution and proving time. The regressions are caused by a 2% increase in executed instructions and 17% increase in paging cycles on both zkVMs.
@@
Unrolling Factor x86 SP1 RISC Zero
#instr. prove exec #instr prove exec
4 +28.1% -50% +24.3% +6.6% -50% +51.4% +39.5%
16 +31.5% -63% +26.4% +1.5% -62% +51.6% +52.7%
When used with -O3 and a high unroll-threshold (aggressive unrolling), loop unrolling becomes beneficial compared with -O3 and no loop unrolling. Specifically, on RISC Zero, execution improves by 16% and proving by 3.7%; on SP1, by 0.5% and 1.7%. In both zkVMs, these performance gains are attributed to by 4–5% decrease in cycle count (explained by a 5% decrease in dynamic instruction count). However, these gains are still modest compared to x86’s $32\times$ speedup improvement. This trend also appears in polybench-gemm, which is also a matrix multiplication benchmark. By increasing the unrolling threshold, we save around 43–44% in terms of cycles (compared to -O3), by executing 44% fewer instructions. This leads to a 11.5% improvement in execution time on RISC Zero and 2% improvement on SP1.
Principle 3 (P3): Performance gains are observed only when loop unrolling leads to a reduction in the executed instructions. Since zkVMs cannot exploit instruction-level parallelism, out-of-order, or superscalar execution, performance gains when unrolling loops depend entirely on reducing instruction count. Compilers targeting zkVMs should prioritize instruction reduction over traditional hardware assumptions.
To validate this principle, we manually unroll loops in RISC-V assembly of program shown in Figure 12, avoiding interference from other passes or vectorization. Table 3 reports speedups for $4\times$ and $16\times$ loop unrolls. This manual effort yields substantial performance gains for both zkVMs, with RISC Zero even surpassing the gains seen on x86. These improvements come mainly from executing fewer loop bookkeeping instructions, such as counter increments, comparisons, and branch jumps. As dynamic instruction count decreases, paging cycles also decrease, by about 40% on average.
@@
Substantial improvement on x86; minimal effect on zkVMs. On polybench-nussinov, simplifycfg improves x86 by 23.6%, but SP1 execution by only 2.6%. Figure 13 shows a minimized function from this benchmark that computes the absolute value of an integer using a branch. simplifycfg replaces the branch with equivalent arithmetic operations (Figures 13(b) and 13(c)). When tested with an array of integers chosen at random, the optimized version runs $2.2\times$ faster on x86, presumably thanks to less branch misprediction overhead. However, on zkVMs, the optimized code increases cycle count by 17.7% on RISC Zero and 7.6% on SP1. These are primarily explained by increases in the number of executed instructions (8%) and paging cycles (17%). As a result, proving time degrades by 18.1% and 7% on RISC Zero and SP1, respectively, while execution time worsens by 8.5% and 7.7%.
Principle 4 (P4): Branches should not always be eliminated, as branch elimination may increase the number of executed instructions by requiring both paths to be evaluated . Compilers targeting zkVMs should be more conservative when removing branches, for two reasons. First, branches are relatively cheap and incur no misprediction penalty compared to traditional CPUs. Second, predicated execution increases proving cost, since both paths must be proven and executed.
To validate this principle, we modify the unoptimized binary of polybench-nussinov by replacing the implementation of the max function with a branchless version implemented with right-shifts and bitwise AND operations. While this change yields a 4% performance improvement on x86, it degrades performance on RISC Zero by 4.4% and on SP1 by 0.5%. This regression is caused by a 3% increase in dynamic instruction count on both zkVMs.
@@ 6. Implications
We demonstrate the practical value of our study’s results by refining existing LLVM passes to improve zkVM performance and then extract several key lessons from our findings.
@@
Change set 1: Adopting a zkVM-specific cost model. The default RISC-V cost model in LLVM is not zkVM-aware, yet instruction costs on zkVMs differ greatly from traditional CPUs. For example, multiplication costs the same as addition, and division is not expensive. Memory access is also far more expensive when it triggers page-ins or page-outs (see Principle 1, Section 5). We address these by updating the RISCVTTIImpl and RISCVTargetLowering classes, which provide target-specific cost models to LLVM’s middle-end so that passes can make better decisions for this architecture.
Change set 2: Updating defaults and heuristics. Many LLVM passes employ heuristics that are tuned for traditional architectures. This tuning may not be suitable for zkVMs. Adjusting these for zkVM-specific instruction costs, such as applying function inlining (Principle 2) or loop unrolling (Principle 3) only when it reduces instructions count can improve the performance of the generated code.
We increase the default inlining threshold and adjust parameters such as inline-call-penalty accordingly. We also update simplifycfg to make branch elimination more conservative and avoiding evaluation of both sides of a predicated operation (Principle 4).
Change set 3: Disabling irrelevant passes. Our RQ1 analysis (Section 4.1) shows that some passes offer little or no benefit on zkVMs, mainly because they rely on hardware features that are unsupported in zkVMs. For example, speculative-execution hoists side-effect-free instructions to reduce branch-misprediction latency on out-of-order, superscalar CPUs. This advantage is irrelevant to zkVMs, which execute RISC-V code strictly in order inside an arithmetic circuit. Similarly, we also disable loop-data-prefetch, and hot-cold-splitting, as they provide no measurable gain.
@@
Results. We evaluate -O3 in our modified LLVM against -O3 in the original toolchain. Figure 14 shows that our LLVM changes improve zkVM performance across a wide range of benchmarks, with execution-time gains exceeding 10% in many cases and reaching up to 45% on RISC Zero for fibonacci. This improvement directly validates our cost model and core principle (Section 5.1): reducing dynamic instruction count translates to proportional performance gains. The fibonacci benchmark experiences a 50% reduction in dynamic instruction count, which corresponds to the 45% execution time improvement on RISC Zero. Specifically, our division cost-model update enables LLVM to select a single remu instruction instead of a series of bitwise operations for remainder calculation.
The modified LLVM outperforms vanilla -O3 on execution time in 39/58 benchmarks on RISC Zero (avg. +4.6%) and 19/58 on SP1 (avg. +1%), with only two and one regressions, respectively. On SP1, proving time improves for 27/58 programs by more than 1%, with only three regressions. On RISC Zero, most proving-time effects are small, with 8 programs improving by more than 1% and 4 programs showing slight regressions. One outlier is regex-match on SP1, where our modifications lead to a 27.3% regression in proving time. This is caused by our changes in the inline’s heuristics, which make SP1 produce 20 proof shards instead of 16. Proof sharding divides the program into smaller proving units; more shards may increase proving time due to additional aggregation overhead.
Notably, our performance improvements are mainly explained by dynamic instruction count reduction: our modifications reduce the number of the executed instructions by 2.6% on average across both zkVMs, with only three benchmarks showing slight increases (max +0.5%).